OpenStudy (anonymous):

Let f be the function (square root(x^4-16x^2)) find the slope of the line normal to the graph of f at x=6.

OpenStudy (amistre64):

the slope is the derivative; what do we find as the derivative?

OpenStudy (anonymous):

i already got the deriv it wud be (2x^2-16x)/ square root of (x^4-16x^2) i just dont know what it means when they ask lline normal to i didnt learn that yet

OpenStudy (anonymous):

that is the deriv right? ><

OpenStudy (amistre64):

4x^3 - 32x --------------- is what i get; then when x=6 ..... 2sqrt(x^4-16x^2) ohh, i read slope of the line; the normal is the perp slope; 1/f'(x) in this case

OpenStudy (amistre64):

in other words; take f'(6) and flip it

OpenStudy (amistre64):

if a tangent slope is: 4/5 the normal slope for that would be : 5/4

OpenStudy (anonymous):

okay hang on gona do that ^^ nd thank you for helping

OpenStudy (amistre64):

forgot a negative

OpenStudy (amistre64):

if a tangent slope is: 4/5 the normal slope for that would be : -5/4 thats better

OpenStudy (anonymous):

oh okay so its kinda like when your doing slopes perpendicular to?

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

not kinda; its the exact same thing :)

OpenStudy (anonymous):

oh they're just out to get us with different wording >< GAH okay so i know how to do it thank you so much ^^

OpenStudy (amistre64):

youre welcome :)

OpenStudy (anonymous):

^^

OpenStudy (anonymous):

uhm sorry again, but i got a weird fraction >< idk if my answer is correct can you just check it please?

OpenStudy (amistre64):

4x^3 - 32x --------------- 2sqrt(x^4-16x^2) 2x^3 - 16x ------------------ sqrt(x^2)srt(x^2-16) 2x^3 - 16x ----------- xsqrt(x^2-16) 2x^2 - 16 ----------- ; x=6 we get .... 56/sqrt(20) sqrt(x^2-16) 56/2sqrt(5) is the slope of the tangent -sqrt(5)/28 should be the slope of the normal

OpenStudy (amistre64):

hmmm, wolf says i mighta had an error

OpenStudy (anonymous):

uhm.. hang on checking with mine

OpenStudy (amistre64):

2(36)-16 = 72-16 = 62-6 = 56 sqrt(6^2-16) = sqrt(36-16) = sqrt(20) = 2sqrt(5) 56/2 = 28 28/sqrt(5) is what i get for f'(6) still

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=derivative+sqrt%28x%5E4-16x%5E2%29%3B+x%3D6 now it likes it ....

OpenStudy (amistre64):

yeah, im good again lol

OpenStudy (anonymous):

wait i think i might have accidently squared a number when i shudnt have /fail hang on again >< really sorry T.T

OpenStudy (anonymous):

for my f'(x) i got 2x^3-16x / square root (x^4-16x)

OpenStudy (anonymous):

i dont get how u got 4x^3 - 32x --------------- 2sqrt(x^4-16x^2)

OpenStudy (anonymous):